# Hölder regularity of the solutions of Fredholm integral equations on upper Ahlfors regular sets

### v3

@@ 1 Introduction
In case $X=Y$, we just say that $Y$ is upper $\upsilon_{Y}$-Ahlfors regular and this is the assumption that has been considered by García-Cuerva and Gatto [8], [9], Gatto [10], [11] in case $X=Y$ in the context of Hölder spaces. See also Edmunds, Kokilashvili and Meskhi [4, Chap. 6] in the frame of Lebsgue spaces.
@@
${\mathbb{D}}_{Y\times Y}\equiv\{(y_{1},y_{y})\in Y\times Y:\,y_{1}=y_{2}\}\,$,
@@ 3 Special classes of potential type kernels in metric spaces
Let $Y$ be a subset of a metric space $(M,d)$. Let $\nu$ be as in (1.1). If $Y$ is bounded, then $\nu(Y)<+\infty$. Indeed, if $Y=\emptyset$, then $\nu(Y)=0$. If instead $Y\neq\emptyset$, then there exists at least an element $\tilde{y}\in Y$ and the assumption that $Y$ is bounded implies that there exists $r\in]0,+\infty[$ such that $Y\subseteq B(\tilde{y},r)$ and accordingly condition (1.1) implies that $\nu(Y)$ is finite.
@@ 4 Composite kernels on upper Ahlfors regular sets
Let $Y$ be a subset of a metric space $(M,d)$. Let $\nu$ be as in (1.1). Let $\nu(Y)<+\infty$. Let $\upsilon_{Y}\in]0,+\infty[$. Let $Y$ be upper $\upsilon_{Y}$-Ahlfors regular with respect to $Y$. Let $s_{1}$, $s_{2}\in[0,\upsilon_{Y}[$. If $s_{1}+s_{2}=\upsilon_{Y}$, then we further assume that $Y$ is strongly upper $\upsilon_{Y}$-Ahlfors regular with respect to $Y$.
@@
Let $Y$ be a subset of a metric space $(M,d)$. Let $\nu$ be as in (1.1). Let $\nu(Y)<+\infty$. Let $\upsilon_{Y}\in]0,+\infty[$. Let $Y$ be upper $\upsilon_{Y}$-Ahlfors regular with respect to $Y$. Let $t_{1}$, $t_{2}\in]0,\upsilon_{Y}]$. Let $K_{l}\in{\mathcal{K}}_{\upsilon_{Y}-t_{l},Y\times Y}$ for $l\in\{1,2\}$. If $t_{1}+t_{2}<\upsilon_{Y}$, then the composite kernel $K_{3}$ of $K_{1}$ and $K_{2}$ is continuous in $Y^{2}\setminus{\mathbb{D}}_{Y\times Y}$ and satisfies the following inequality
@@
Let $Y$ be a bounded subset of a metric space $(M,d)$. Let $\nu$ be as in (1.1). Let $\upsilon_{Y}\in]0,+\infty[$. Let $s\in[0,\upsilon_{Y}[$. Let $Y$ be upper $\upsilon_{Y}$-Ahlfors regular with respect to $Y$. Let $K\in{\mathcal{K}}_{s,Y\times Y}$.
@@
Proof. By Remark 3.4, we have $\nu(Y)<+\infty$. If $s=0$, then Theorem 4.16 (ii) implies that $K^{(2)}$ admits a continuous and bounded extension to $Y^{2}$. Thus we can assume that $s\in]0,\upsilon_{Y}[$. By assumption,
@@
Let $Y$ be a subset of a metric space $(M,d)$. Let $\nu$ be as in (1.1). Let $\upsilon_{Y}\in]0,+\infty[$. Let $Y$ be upper $\upsilon_{Y}$-Ahlfors regular with respect to $X$. Let $\nu(Y)<+\infty$ whenever $r_{X,Y,\upsilon_{Y}}<+\infty$. Let $s_{1}$, $s_{2}\in[0,\upsilon_{Y}[$, $a\in]0,+\infty[$. Then there exists $c_{s_{1},s_{2},a,Y}\in]0,+\infty[$ such that
@@
and accordingly that $d(x^{\prime},x^{\prime\prime})\leq d(\xi,y)$ and thus again a contradiction. By inequality (4.26) and equality (4.27), we obtain that
@@
for all $x^{\prime}$, $^{\prime\prime}\in Y, xx$$^{\prime}\neq x^{\prime\prime}$, $\xi\in\{x^{\prime},x^{\prime\prime}\}$, $y\in Y\setminus B(x^{\prime},2d(x^{\prime},x^{\prime\prime}))$. Next we note that again Lemma 3.5 (iii) implies that
@@
$\qquad\leq 2^{s_{1}}d(\xi,y)^{-s_{1}}\int_{Y\cap B(\xi,ad(x^{\prime},x^{\prime\prime})}\frac{d\nu(\eta)}{d(\eta,y)^{s_{2}}}$
$\qquad\leq 2^{s_{1}}d(\xi,y)^{-s_{1}}c^{\prime\prime}_{s_{2},Y,Y}a^{\upsilon_{Y}-s_{2}}d(x^{\prime},x^{\prime\prime})^{\upsilon_{Y}-s_{2}}$
@@
Let $Y$ be a bounded subset of a metric space $(M,d)$. Let $\nu$ be as in (1.1). Let $\upsilon_{Y}\in]0,+\infty[$. Let $Y$ be upper $\upsilon_{Y}$-Ahlfors regular with respect to $Y$. Let $s_{1}$, $t_{1}\in[0,\upsilon_{Y}[$. Let $s_{3}\in[0,+\infty[$. Let $s_{2}^{\prime}\in[0,\upsilon_{Y}[$, $s_{2}^{\prime\prime}\in[0,s_{3}]$, $s_{2}\equiv s_{2}^{\prime}+s_{2}^{\prime\prime}$. If either $s_{1}+t_{1}=\upsilon_{Y}$ or $s_{2}^{\prime}+t_{1}=\upsilon_{Y}$, then we further assume that $Y$ is strongly upper $\upsilon_{Y}$-Ahlfors regular with respect to $Y$. Let $K_{1}\in{\mathcal{K}}_{s_{1},s_{2},s_{3}}(Y\times Y)$, $K_{2}\in{\mathcal{K}}_{t_{1},Y\times Y}$. Let $K_{3}$ be the composite kernel of $K_{1}$ and $K_{2}$ (cf. (4.17)). Then the following statements hold.
@@
Proof. By Remark 3.4, we have $\nu(Y)<+\infty$. In all cases (i)–(ix), the membership of $K_{3}$ in the class ${\mathcal{K}}_{s,Y\times Y}$ for the choice of $s$ as in the first index is an immediate consequence of Theorem 4.16. Then we note that if $x^{\prime}$, $^{\prime\prime}\in Y, xx$$^{\prime}\neq x^{\prime\prime}$, $y\in Y\setminus B(x^{\prime},2d(x^{\prime},x^{\prime\prime}))$, then the triangular inequality, Proposition 4.1, Lemma 4.24, the elementary inclusion $B(x^{\prime},2d(x^{\prime},x^{\prime\prime}))\subseteq B(x^{\prime\prime},3d(x^{\prime},x^{\prime\prime}))$ and the elementary inequality $d(y,x^{\prime})\geq 2d(x^{\prime},x^{\prime\prime})$ that holds whenever $y\in Y\setminus B(x^{\prime},2d(x^{\prime},x^{\prime\prime}))$ imply that
@@
$\quad\leq 2^{-s_{2}^{\prime\prime}}d(x^{\prime},x^{\prime\prime})^{s_{3}-s_{2}^{\prime\prime}}\|K\|_{{\mathcal{K}}_{s_{1},s_{2},s_{3}}(Y\times Y)}\|K_{2}\|_{{\mathcal{K}}_{t_{1},Y\times Y}}c_{X,Y,Z}^{(s_{2}^{\prime},t_{1})}I(x^{\prime},y)$
@@ 5 A regularity theorem for Fredholm integral equations of the second kind on upper Ahlfors regular sets
Let $Y$ be a bounded subset of a metric space $(M,d)$. Let $\nu$ be as in (1.1). Let $\upsilon_{Y}\in]0,+\infty[$. Let $s\in[0,\upsilon_{Y}[$. Let $Y$ be upper $\upsilon_{Y}$-Ahlfors regular with respect to $Y$. Let $K\in{\mathcal{K}}_{s,Y\times Y}$. Let $A[K,\cdot]$ be the integral operator in $L^{2}_{\nu}(Y)$ that is associated to the kernel $K$ (cf. (1.5)). If $g\in C^{0}_{b}(Y)$ and if $\mu\in L^{2}_{\nu}(Y)$ satisfies the inequality
$\mu-A[K,\mu]=g\qquad\nu-{\mathrm{a.e.\ in}}\ Y\,.$ (5.2)
@@
and thus equality equality (5.3) holds for $r+1$. Then the induction principle implies that equality (5.3) holds for all $r\in{\mathbb{N}}\setminus\{0\}$. By Remark 3.4, we have $\nu(Y)<+\infty$. Since $s\in[0,\upsilon_{Y}[$, and $K$ is continuous outside of the diagonal, it is known that the operator $A[\cdot]$ is linear and continuous from $L^{\infty}_{\nu}(Y)$ to $C^{0}_{b}(Y)$ (cf. e.g., [16, Prop. 4.1], [17, Prop. 4.3]). Then our assumption that $g\in L^{\infty}_{\nu}(Y)$, implies that $A[g]\in C^{0}_{b}(Y)$ and similarly that $A^{j}[g]\in C^{0}_{b}(Y)$ for all $j\in{\mathbb{N}}$. Since $Y$ is bounded and the kernel $K$ has a weak singularity and is continuous on $Y^{2}\setminus{\mathbb{D}}_{Y\times Y}$, Corollary 4.23 ensures that there exists $r\in{\mathbb{N}}\setminus\{0\}$ such that $K^{(r)}$ has a continuous and bounded extension to $Y^{2}$. Then the continuity and boundedness of the extension of $K^{(r)}$ to $Y^{2}$ and classical theorems of continuity for integrals depending on a parameter imply that $A^{r}[\mu]\in C^{0}_{b}(Y)$ (cf. e.g., [16, Prop. 4.1], [17, Prop. 4.3]). Then we have
@@
Let $Y$ be a bounded subset of a metric space $(M,d)$. Let $\nu$ be as in (1.1). Let $\upsilon_{Y}\in]0,+\infty[$. Let $Y$ be upper $\upsilon_{Y}$-Ahlfors regular with respect to $Y$. Let $\theta\in]0,1]$,
@@
Let $K\in{\mathcal{K}}_{s_{1},s_{2},s_{3}}(X\times Y)$. Let $A[K,\cdot]$ be the integral operator in $L^{2}_{\nu}(Y)$ that is associated to the kernel $K$ (cf. (1.5)).
@@
Proof. By Theorem 5.1, we know that $\mu\in C_{b}^{0}(Y)$. By Remark 3.4, we have $\nu(Y)<+\infty$. Then under our assumptions, [16, Prop. 5.2] ensures that $A[K,\cdot]$ is linear and continuous from $C_{b}^{0}(Y)$ to $C_{b}^{0,\varpi(\cdot)}(Y)$. Hence, $A[K,\mu]$ belongs to $C_{b}^{0,\varpi(\cdot)}(Y)$. Since $g\in C^{0,\theta}_{b}(Y)$, equality (6.3) implies that $\mu\in C_{b}^{0,\max\{r^{\theta},\varpi(r)\}}(Y)$. $\Box$
@@
Let $A[K,\cdot]$ be the integral operator in $L^{2}_{\nu}(\partial\Omega)$ that is associated to the kernel $K$ (cf. (1.5) with $Y=\partial\Omega$). If $g\in C^{0,\alpha}_{b}(\partial\Omega)$ and $\mu\in L^{2}_{\nu}(\partial\Omega)$ satisfy the equality
@@
We note that under the assumptions of the previous example, if $\Omega$ is of class $C^{1,\alpha}$, then the kernel of the double layer potential associated to a fundamental solution of a second order elliptic operator with constant coefficients is known to be of class ${\mathcal{K}}_{(n-1)-\alpha,n-\alpha,1}((\partial\Omega)\times(\partial\Omega))$ (see reference [3, Rem. 6.1] with Dondi).
@@
Let $Y$ be a bounded subset of a metric space $(M,d)$. Let $\nu$ be as in (1.1). Let $\upsilon_{Y}\in]0,+\infty[$. Let $\theta\in]0,1]$. Let
@@
Let $Y$ be upper $\upsilon_{Y}$-Ahlfors regular with respect to $Y$. If $s_{2}-\beta=\upsilon_{Y}$, we further require that $Y$ be strongly upper $\upsilon_{Y}$-Ahlfors regular with respect to $X$.
@@
Let $K\in{\mathcal{K}}_{s_{1},s_{2},s_{3}}(X\times Y)$. Let $A[K,\cdot]$ be the integral operator in $L^{2}_{\nu}(Y)$ that is associated to the kernel $K$ (cf. (1.5)). Let
@@
Proof. By Theorem 5.1, we know that $\mu\in C_{b}^{0}(Y)$. By assumption (6.8), we have $A[K,1]\in C^{0,\omega(\cdot)}_{b}(Y)$. By Remark 3.4, we have $\nu(Y)<+\infty$. Then [16, Prop. 5.11] and Remark 2.2 imply that $A[K,\mu]\in C_{b}^{0,\omega(\cdot)}(Y)$. Since $g\in C^{0,\theta}_{b}(Y)$, equality (6.9) implies that $\mu\in C_{b}^{0,\max\{r^{\theta},\omega(r)\}}(Y)$. $\Box$
We note that under the assumptions of Theorem 6.6, if $C_{b}^{0,\omega(\cdot)}(Y)\subseteq C_{b}^{0,\theta}(Y)$, then we can infer that a solution $\mu\in L^{2}_{\nu}(Y)$ of the Fredholm integral equation (6.9) is actually of class $C_{b}^{0,\theta}(Y)$ as the datum $g$.

### v4

@@ 1 Introduction
In case $X=Y$, we just say that $Y$ is upper $\upsilon_{Y}$-Ahlfors regular and this is the assumption that has been considered by García-Cuerva and Gatto [8], [9], Gatto [10], [11] in the context of Hölder spaces. See also Edmunds, Kokilashvili and Meskhi [4, Chap. 6] in the frame of Lebsgue spaces.
@@
${\mathbb{D}}_{Y\times Y}\equiv\{(y_{1},y_{2})\in Y\times Y:\,y_{1}=y_{2}\}\,$,
@@ 3 Special classes of potential type kernels in metric spaces
Let $X$ and $Y$ be subsets of a metric space $(M,d)$. Let $\nu$ be as in (1.1) . If $Y\neq\emptyset$, then we further require that $X\neq\emptyset$. If $Y$ is bounded, then $\nu(Y)<+\infty$. Indeed, if $Y=\emptyset$, then $\nu(Y)=0$. If instead $Y\neq\emptyset$, then there exist at least an element $\tilde{x}\in X$ and an element $\tilde{y}\in Y$ and the assumption that $Y$ is bounded together with the triangular inequality imply that there exists $r\in]0,+\infty[$ such that $Y\subseteq B(\tilde{x},r)$ and accordingly condition (1.1) implies that $\nu(Y)$ is finite.
@@ 4 Composite kernels on upper Ahlfors regular sets
Let $Y$ be a subset of a metric space $(M,d)$. Let $\nu$ be as in (1.1) with $X=Y$. Let $\nu(Y)<+\infty$. Let $\upsilon_{Y}\in]0,+\infty[$. Let $Y$ be upper $\upsilon_{Y}$-Ahlfors regular with respect to $Y$. Let $s_{1}$, $s_{2}\in[0,\upsilon_{Y}[$. If $s_{1}+s_{2}=\upsilon_{Y}$, then we further assume that $Y$ is strongly upper $\upsilon_{Y}$-Ahlfors regular with respect to $Y$.
@@
Let $Y$ be a subset of a metric space $(M,d)$. Let $\nu$ be as in (1.1) with $X=Y$. Let $\nu(Y)<+\infty$. Let $\upsilon_{Y}\in]0,+\infty[$. Let $Y$ be upper $\upsilon_{Y}$-Ahlfors regular with respect to $Y$. Let $t_{1}$, $t_{2}\in]0,\upsilon_{Y}]$. Let $K_{l}\in{\mathcal{K}}_{\upsilon_{Y}-t_{l},Y\times Y}$ for $l\in\{1,2\}$. If $t_{1}+t_{2}<\upsilon_{Y}$, then the composite kernel $K_{3}$ of $K_{1}$ and $K_{2}$ is continuous in $Y^{2}\setminus{\mathbb{D}}_{Y\times Y}$ and satisfies the following inequality
@@
Let $Y$ be a bounded subset of a metric space $(M,d)$. Let $\nu$ be as in (1.1) with $X=Y$. Let $\upsilon_{Y}\in]0,+\infty[$. Let $s\in[0,\upsilon_{Y}[$. Let $Y$ be upper $\upsilon_{Y}$-Ahlfors regular with respect to $Y$. Let $K\in{\mathcal{K}}_{s,Y\times Y}$.
@@
Proof. By Remark 3.4 with $X=Y$, we have $\nu(Y)<+\infty$. If $s=0$, then Theorem 4.16 (ii) implies that $K^{(2)}$ admits a continuous and bounded extension to $Y^{2}$. Thus we can assume that $s\in]0,\upsilon_{Y}[$. By assumption,
@@
Let $Y$ be a subset of a metric space $(M,d)$. Let $\nu$ be as in (1.1) with $X=Y$. Let $\upsilon_{Y}\in]0,+\infty[$. Let $Y$ be upper $\upsilon_{Y}$-Ahlfors regular with respect to $Y$. Let $\nu(Y)<+\infty$ whenever $r_{Y,Y,\upsilon_{Y}}<+\infty$. Let $s_{1}$, $s_{2}\in[0,\upsilon_{Y}[$, $a\in]0,+\infty[$. Then there exists $c_{s_{1},s_{2},a,Y}\in]0,+\infty[$ such that
@@
and accordingly that $d(x^{\prime},x^{\prime\prime})\leq d(\xi,y)<d(x^{\prime},x^{\prime\prime})$ and thus again a contradiction. By inequality (4.26) and equality (4.27), we obtain that
@@
for all $x^{\prime}$, $^{\prime\prime}\in Y, xx$$^{\prime}\neq x^{\prime\prime}$, $\xi\in\{x^{\prime},x^{\prime\prime}\}$, $y\in Y\setminus B(x^{\prime},2d(x^{\prime},x^{\prime\prime}))$. Next we note that
$A(x^{\prime},x^{\prime\prime},a)\cap\{\eta\in Y:d(\xi,\eta)\geq d(\xi,y)/2\}\subseteq Y\cap B(y,3ad(x^{\prime},x^{\prime\prime}))$
for all $x^{\prime}$, $^{\prime\prime}\in Y, xx$$^{\prime}\neq x^{\prime\prime}$, $\xi\in\{x^{\prime},x^{\prime\prime}\}$, $y\in Y\setminus B(x^{\prime},2d(x^{\prime},x^{\prime\prime}))$. Hence, again Lemma 3.5 (iii) implies that
@@
$\qquad\leq 2^{s_{1}}d(\xi,y)^{-s_{1}}\int_{Y\cap B(y,3ad(x^{\prime},x^{\prime\prime}))}\frac{d\nu(\eta)}{d(\eta,y)^{s_{2}}}$
$\qquad\leq 2^{s_{1}}d(\xi,y)^{-s_{1}}c^{\prime\prime}_{s_{2},Y,Y}(3a)^{\upsilon_{Y}-s_{2}}d(x^{\prime},x^{\prime\prime})^{\upsilon_{Y}-s_{2}}$
@@
Let $Y$ be a bounded subset of a metric space $(M,d)$. Let $\nu$ be as in (1.1) with $X=Y$. Let $\upsilon_{Y}\in]0,+\infty[$. Let $Y$ be upper $\upsilon_{Y}$-Ahlfors regular with respect to $Y$. Let $s_{1}$, $t_{1}\in[0,\upsilon_{Y}[$. Let $s_{3}\in[0,+\infty[$. Let $s_{2}^{\prime}\in[0,\upsilon_{Y}[$, $s_{2}^{\prime\prime}\in[0,s_{3}]$, $s_{2}\equiv s_{2}^{\prime}+s_{2}^{\prime\prime}$. If either $s_{1}+t_{1}=\upsilon_{Y}$ or $s_{2}^{\prime}+t_{1}=\upsilon_{Y}$, then we further assume that $Y$ is strongly upper $\upsilon_{Y}$-Ahlfors regular with respect to $Y$. Let $K_{1}\in{\mathcal{K}}_{s_{1},s_{2},s_{3}}(Y\times Y)$, $K_{2}\in{\mathcal{K}}_{t_{1},Y\times Y}$. Let $K_{3}$ be the composite kernel of $K_{1}$ and $K_{2}$ (cf. (4.17)). Then the following statements hold.
@@
Proof. By Remark 3.4 with $X=Y$, we have $\nu(Y)<+\infty$. In all cases (i)–(ix), the membership of $K_{3}$ in the class ${\mathcal{K}}_{s,Y\times Y}$ for the choice of $s$ as in the first index is an immediate consequence of Theorem 4.16. Then we note that if $x^{\prime}$, $^{\prime\prime}\in Y, xx$$^{\prime}\neq x^{\prime\prime}$, $y\in Y\setminus B(x^{\prime},2d(x^{\prime},x^{\prime\prime}))$, then the triangular inequality, Proposition 4.1, Lemma 4.24, the elementary inclusion $B(x^{\prime},2d(x^{\prime},x^{\prime\prime}))\subseteq B(x^{\prime\prime},3d(x^{\prime},x^{\prime\prime}))$ and the elementary inequality $d(y,x^{\prime})\geq 2d(x^{\prime},x^{\prime\prime})$ that holds whenever $y\in Y\setminus B(x^{\prime},2d(x^{\prime},x^{\prime\prime}))$ imply that
@@
$\quad\leq 2^{-s_{2}^{\prime\prime}}d(x^{\prime},x^{\prime\prime})^{s_{3}-s_{2}^{\prime\prime}}\|K\|_{{\mathcal{K}}_{s_{1},s_{2},s_{3}}(Y\times Y)}\|K_{2}\|_{{\mathcal{K}}_{t_{1},Y\times Y}}c_{Y,Y,Y}^{(s_{2}^{\prime},t_{1})}I(x^{\prime},y)$
@@ 5 A regularity theorem for Fredholm integral equations of the second kind on upper Ahlfors regular sets
Let $Y$ be a bounded subset of a metric space $(M,d)$. Let $\nu$ be as in (1.1) with $X=Y$. Let $\upsilon_{Y}\in]0,+\infty[$. Let $s\in[0,\upsilon_{Y}[$. Let $Y$ be upper $\upsilon_{Y}$-Ahlfors regular with respect to $Y$. Let $K\in{\mathcal{K}}_{s,Y\times Y}$. Let $A[K,\cdot]$ be the integral operator in $L^{2}_{\nu}(Y)$ that is associated to the kernel $K$ (cf. (1.5)). If $g\in C^{0}_{b}(Y)$ and if $\mu\in L^{2}_{\nu}(Y)$ satisfies the inequality
$\mu-A[K,\mu]=g\qquad\nu-{\mathrm{a.e.\ in}}\ Y\,$, (5.2)
@@
and thus equality equality (5.3) holds for $r+1$. Then the induction principle implies that equality (5.3) holds for all $r\in{\mathbb{N}}\setminus\{0\}$. By Remark 3.4 with $X=Y$, we have $\nu(Y)<+\infty$. Since $s\in[0,\upsilon_{Y}[$, and $K$ is continuous outside of the diagonal, it is known that the operator $A[\cdot]$ is linear and continuous from $L^{\infty}_{\nu}(Y)$ to $C^{0}_{b}(Y)$ (cf. e.g., [16, Prop. 4.1], [17, Prop. 4.3]). Then our assumption that $g\in L^{\infty}_{\nu}(Y)$, implies that $A[g]\in C^{0}_{b}(Y)$ and similarly that $A^{j}[g]\in C^{0}_{b}(Y)$ for all $j\in{\mathbb{N}}$. Since $Y$ is bounded and the kernel $K$ has a weak singularity and is continuous on $Y^{2}\setminus{\mathbb{D}}_{Y\times Y}$, Corollary 4.23 ensures that there exists $r\in{\mathbb{N}}\setminus\{0\}$ such that $K^{(r)}$ has a continuous and bounded extension to $Y^{2}$. Then the continuity and boundedness of the extension of $K^{(r)}$ to $Y^{2}$ and classical theorems of continuity for integrals depending on a parameter imply that $A^{r}[\mu]\in C^{0}_{b}(Y)$. Then we have
@@
Let $Y$ be a bounded subset of a metric space $(M,d)$. Let $\nu$ be as in (1.1) with $X=Y$. Let $\upsilon_{Y}\in]0,+\infty[$. Let $Y$ be upper $\upsilon_{Y}$-Ahlfors regular with respect to $Y$. Let $\theta\in]0,1]$,
@@
Let $K\in{\mathcal{K}}_{s_{1},s_{2},s_{3}}(Y\times Y)$. Let $A[K,\cdot]$ be the integral operator in $L^{2}_{\nu}(Y)$ that is associated to the kernel $K$ (cf. (1.5) with $X=Y$).
@@
Proof. By Theorem 5.1, we know that $\mu\in C_{b}^{0}(Y)$. By Remark 3.4 with $X=Y$, we have $\nu(Y)<+\infty$. Then under our assumptions, [16, Prop. 5.2] ensures that $A[K,\cdot]$ is linear and continuous from $C_{b}^{0}(Y)$ to $C_{b}^{0,\varpi(\cdot)}(Y)$. Hence, $A[K,\mu]$ belongs to $C_{b}^{0,\varpi(\cdot)}(Y)$. Since $g\in C^{0,\theta}_{b}(Y)$, equality (6.3) implies that $\mu\in C_{b}^{0,\max\{r^{\theta},\varpi(r)\}}(Y)$. $\Box$
@@
Let $A[K,\cdot]$ be the integral operator in $L^{2}_{\nu}(\partial\Omega)$ that is associated to the kernel $K$ (cf. (1.5) with $X=Y=\partial\Omega$). If $g\in C^{0,\alpha}_{b}(\partial\Omega)$ and $\mu\in L^{2}_{\nu}(\partial\Omega)$ satisfy the equality
@@
We note that under the assumptions of the previous example, if $\Omega$ is of class $C^{1,\alpha}$, then the kernel of the double layer potential that is associated to a fundamental solution of a second order elliptic operator with constant coefficients is known to be of class ${\mathcal{K}}_{(n-1)-\alpha,n-\alpha,1}((\partial\Omega)\times(\partial\Omega))$ (see reference [3, Rem. 6.1] with Dondi).
@@
Let $Y$ be a bounded subset of a metric space $(M,d)$. Let $\nu$ be as in (1.1) with $X=Y$. Let $\upsilon_{Y}\in]0,+\infty[$. Let $\theta\in]0,1]$. Let
@@
Let $Y$ be upper $\upsilon_{Y}$-Ahlfors regular with respect to $Y$. If $s_{2}-\beta=\upsilon_{Y}$, we further require that $Y$ be strongly upper $\upsilon_{Y}$-Ahlfors regular with respect to $Y$.
@@
Let $K\in{\mathcal{K}}_{s_{1},s_{2},s_{3}}(Y\times Y)$. Let $A[K,\cdot]$ be the integral operator in $L^{2}_{\nu}(Y)$ that is associated to the kernel $K$ (cf. (1.5) with $X=Y$). Let
@@
Proof. By assumption, we know that $\mu\in C_{b}^{0,\beta}(Y)$. By assumption (6.8), we have $A[K,1]\in C^{0,\omega(\cdot)}_{b}(Y)$. By Remark 3.4 with $X=Y$, we have $\nu(Y)<+\infty$. Then [16, Prop. 5.11] and Remark 2.2 imply that $A[K,\mu]\in C_{b}^{0,\omega(\cdot)}(Y)$. Since $g\in C^{0,\theta}_{b}(Y)$, equality (6.9) implies that $\mu\in C_{b}^{0,\max\{r^{\theta},\omega(r)\}}(Y)$. $\Box$
We note that under the assumptions of Theorem 6.6, if $C_{b}^{0,\omega(\cdot)}(Y)\subseteq C_{b}^{0,\theta}(Y)$, then we can infer that $\mu$ is actually of class $C_{b}^{0,\theta}(Y)$ as the datum $g$.